Write the standard form of the equation of the line that passes through the point (6,4) and has a slope of 8/5.
Answer by yourstruly(1)

First you should remember the formula: y-y1= m(x-x1)
m= gradient/slope
x1,y1= coordinates of the point that the line passes through
In this question our slope is 5/8 and our x and y coordinates are 6 and 4 respectively.
When we plug this into the formula above we get:
y-4= 5/8(x-6)
y-4= 5/8x-15/4
y=5/8x+1/4
Now, we have to put this equation into standard form which is
ax+by+c=0
8y=5x+2
5x+2-8y=0----- this is the answer